Reduced new FF matrix acquire with cap: TN 20110826_235832

test fiber with W: ok

RR put between 200and400um

Slew test @12.5

Act 327DSP jumped. Removed from ICC.

Restarted @ 13:00

Act 568DSP humped to 2 m. -> removed.

Restarted @ 13.12--->vulcano effect on 568

DB releasing, testing where 568 gives good readings and tight.

Restarted test@ 13:31

seen bad behavior around 30 single spot

Test end @13.41

-- MarcoXompero - 27 Aug 2011
Topic revision: r2 - 27 Aug 2011, MarcoXompero
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback